OpenStudy (anonymous):

Select the ordered pair from the choices below that is a solution to the following system of equations: 3x = y + 18 y + 7 = -30x (7, 3) (-2, 12) (1/3, -17) (-4, 2)

OpenStudy (anonymous):

You can solve this system of equations with substitution or subtraction. Do you have a preference of method?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I would recommend substitution. It works well for all systems of equations. Start by picking one equation and solve for one variable. My suggestion is to solve either equation for y since that will avoid fractions.

OpenStudy (anonymous):

3x = y + 18 y + 7 = -30x i.e. y= -30x-7 i.e. 3x= -30x-7+18 33x=11 x=11/33=1/3 y=-10-7=-17 thus (x,y)=(1/3, -17) is the correct ordered pair

OpenStudy (anonymous):

thank you